OpenStudy (anonymous):

Paulo is installing a circular pond in his back yard. He is surrounding the pond with decorative bricks. The pond is 5 feet in diameter. What is the length of the decorative brick border that Paulo needs to surround the pond? Use twenty-two sevenths for pi. seven and six-sevenths feet fifteen and five-sevenths feet nineteen and nine-fourteenths feet thirty-one and three-sevenths feet @cp9454

OpenStudy (anonymous):

@math&ing001

OpenStudy (math&ing001):

You'll just need to find the diameter of the circle, do you know how to do that ?

OpenStudy (anonymous):

no

OpenStudy (math&ing001):

Sorry I meant the perimeter. Perimeter = diameter * pi = 5 * 22/7

OpenStudy (anonymous):

3

OpenStudy (math&ing001):

It's not option 3, perimeter = 5 * 22/7 = 110/7 = 15 + 5/7. Option B.